The two stones were thrown vertically downwards from a height of 50 m towards the surface of the lake, with the distance between their throws being 1 s. Both stones hit the surface of the lake at the same time. The initial speed of the first stone is 2m / s. Determine the speed at which the second stone is thrown vertically downwards.



1
Expert's answer
2021-03-04T11:53:49-0500
"h=v_1t+0.5gt^2\\50=2t+0.5(9.8)t^2\\t=3\ s"

"h=v_2(t-1)+0.5g(t-1)^2\\50=v_2(3-1)+0.5(9.8)(3-1)^2\\v_2=15.2\frac{m}{s}"